# https://sites.google.com/view/echoscene

### v1

@@ Abstract
We present EchoScene, an interactive and controllable generative model that generates 3D indoor scenes on scene graphs. EchoScene leverages a dual-branch diffusion model that dynamically adapts to scene graphs. Existing methods struggle to handle scene graphs due to varying numbers of nodes, multiple edge combinations, and manipulator-induced node-edge operations. EchoScene overcomes this by associating each node with a denoising process and enables collaborative information exchange, enhancing controllable and consistent generation aware of global constraints. This is achieved through an information echo scheme in both shape and layout branches. At every denoising step, all processes share their denoising data with an information exchange unit that combines these updates using graph convolution. The scheme ensures that the denoising processes are influenced by a holistic understanding of the scene graph, facilitating the generation of globally coherent scenes. The resulting scenes can be manipulated during inference by editing the input scene graph and sampling the noise in the diffusion model. Extensive experiments validate our approach, which maintains scene controllability and surpasses previous methods in generation fidelity. Moreover, the generated scenes are of high quality and thus directly compatible with off-the-shelf texture generation. Code and trained models are open-sourced.
@@ 2 Related work
Semantic Scene Graph. Scene graphs have become a fundamental tool for semantic scene understanding, offering a structured and symbolic representation of scenes through nodes (objects) and edges (relationships) [21]. Their versatility spans various domains, including image manipulation, caption generation, and visual question answering, demonstrating their capacity to enrich 2D image generation [20, 14, 25]. Beyond 2D, scene graphs have been extended to 3D scene understanding [54, 61], dynamic modeling [45], and notably, controllable scene synthesis [27], showcasing their adaptability in representing complex spatial and semantic relations in multi-dimensional spaces [54, 45, 27, 60, 38, 57]. EchoScene proposes an indoor 3D scene generation method conditioned fully on semantic scene graphs coupled with a scene controlling ability.
Diffusion Models. Applications of diffusion models [48] have witnessed rapid expansion, from generating intricate images to modeling complex distributions in various data types [19, 50]. Recent advances have focused on improving their flexibility and realism, with significant efforts to enhance conditional generation and refine the models’ understanding of input conditions [37, 58, 8, 4]. Score Based Distillation (SDS) was proposed to enable training Neural Radiance Fields through pretrained diffusion models [42]. A line of work addressed text-to-3D via SDS[46, 32], studied the compositionality of scenes via SDS[31, 13, 17, 41], and more recently via Gaussian Splatting [59]. These methods can create realistic 3D assets from text via a small number of images; however, they are limited in terms of number of objects within a scene while capturing their relationships. Comparably, a set of works used diffusion models to learn 3D shapes explicitly [12, 33]. Finally, diffusion models have recently been shown to be useful for graph generative processes [53, 52, 24]. EchoScene draws from the recent advances in diffusion models, where both shape and layout generation are through graph diffusion processes.
@@ 5.1 Scene Fidelity
5.2 Graph Constraints
We optionally manipulate scene graphs in the latent by either adding a node and relevant edges to the graph or changing relations between two nodes, as illustrated in Fig. 2. As shown in the top and middle row of Tab. 2, for both node addition and relation change, EchoLayout and EchoScene perform better on most metrics. 3D-SLN, CommonScenes, and Graph-to-3D, for instance, loose the constraints smaller/bigger than and close by after graph manipulation, while ours largely keeps it. We believe that the improvement is caused by our scene graph diffusion strategy, where a graph-consistent sample is generated within a realistic appearance distribution. The decrease in symmertical category compared with CommonScenes is likely caused by the same effect. Inspection of the data reveals that symmertical is an extremely rare label in the supplementary of CommonScenes [56], with an occurrence of only 0.9%. Thus, the learned part of the distribution is underrepresented and less accurate. The residual echo correction diffuses the latent code towards the learned distribution, resulting in a slight variation. In the bottom row, we report the results without manipulation. All baselines show good performance, yet our EchoLayout and EchoScene are still competitive on their own track. Notably, even though InstructScene shows impressive performance on the global generation fidelity shown in Tab. 1, it struggles to maintain the multiple graph constraints simultaneously, as we described in Sec. 1, while our information echo scheme can sufficiently aggregate relations and condition each process to achieve better results. Distribution-dependant effects from data equally manifest in the results where we can see some drop for symmertical again. Interestingly, the effect is less pronounced for a lower parametric final task such as layout generation. This can be explained by the fact that the variability of bounding boxes is smaller compared to shape variations.
@@ 5.3 Ablation study
We ablate three parts of the full version and report the performance in Tab. 4. Firstly, we ablate the time embeddings $\pi(t)$ to check the influence on the model without explicit time information. We observe some marginal decrease compared with the full version, indicating even without $\pi(t)$, the model still learns temporal information through the denoising data at every step. Secondly, we exclude the shape echoes to degrade the shape branch as the one in CommonScenes. We clearly see a drop in the scene fidelity evaluation, where ‘mSG’ focusing on the layouts is not influenced. Thirdly, we substitute cross-attention for concatenation to check if the way of the condition influences the information echo scheme. We also observe a drop in all metrics, especially for fidelity.
@@ 5.4 Application
EchoScene can serve multiple purposes, including but not limited to robotic imagination and manipulation [55]. However, the limitation of EchoScene is that it can generate semantic yet only textureless scenes, which prohibits it from downstream tasks requiring photorealistic textures. Yet, the generated scenes are of high quality, so that the limitation can be alleviated by directly equipping EchoScene with a subsequent off-the-shelf texture generator. Here, we demonstrate that textured scenes can be generated via SceneTex [9] in Fig. 5.
@@ Appendix 0.E Additional Experimental Details
Graph-to-3D series [15]. This series includes one generative method and three object retrieval methods. First, the full generative version Graph-to-3D, stacking two VAE-based branches for object and layout generation, respectively. Second, Graph-to-Box, the single layout branch focusing on the object retrieval task. Third, Progressive, a modified baseline upon Graph-to-Box, specifically adding objects one by one in an autoregressive manner. Fourth, 3D-SLN [34], sharing the same architecture as Graph-to-Box, but without layout standardization during training. We follow the illustration in the supplementary of CommonScenes [56].
CommonScenes series [57]. This series includes the fully generative version CommonScenes, and its layout branch for object retrieval, CommonLayout. We follow the illustration in the supplementary of CommonScenes [56].

### v2

# 3Munich Center for Machine Learning    4Google

@@ Abstract
We present EchoScene, an interactive and controllable generative model that generates 3D indoor scenes on scene graphs. EchoScene leverages a dual-branch diffusion model that dynamically adapts to scene graphs. Existing methods struggle to handle scene graphs due to varying numbers of nodes, multiple edge combinations, and manipulator-induced node-edge operations. EchoScene overcomes this by associating each node with a denoising process and enables collaborative information exchange, enhancing controllable and consistent generation aware of global constraints. This is achieved through an information echo scheme in both shape and layout branches. At every denoising step, all processes share their denoising data with an information exchange unit that combines these updates using graph convolution. The scheme ensures that the denoising processes are influenced by a holistic understanding of the scene graph, facilitating the generation of globally coherent scenes. The resulting scenes can be manipulated during inference by editing the input scene graph and sampling the noise in the diffusion model. Extensive experiments validate our approach, which maintains scene controllability and surpasses previous methods in generation fidelity. Moreover, the generated scenes are of high quality and thus directly compatible with off-the-shelf texture generation. Our code and models are open-sourced.
@@ 2 Related work
Semantic Scene Graph. Scene graphs have become a fundamental tool for semantic scene understanding, offering a structured and symbolic representation of scenes through nodes (objects) and edges (relationships) [21]. Their versatility spans various domains, including image manipulation, caption generation, and visual question answering, demonstrating their capacity to enrich 2D image generation [20, 14, 27]. Beyond 2D, scene graphs have been extended to 3D scene understanding [56, 62], dynamic modeling [47], and notably, controllable scene synthesis [29], showcasing their adaptability in representing complex spatial and semantic relations in multi-dimensional spaces [56, 47, 29, 61, 40, 58, 26]. EchoScene proposes an indoor 3D scene generation method conditioned fully on semantic scene graphs coupled with a scene controlling ability.
Diffusion Models. Applications of diffusion models [50] have witnessed rapid expansion, from generating intricate images to modeling complex distributions in various data types [19, 52]. Recent advances have focused on improving their flexibility and realism, with significant efforts to enhance conditional generation and refine the models’ understanding of input conditions [39, 59, 25, 8, 4]. Score Distillation Sampling (SDS) was proposed to enable training Neural Radiance Fields through pretrained diffusion models [44]. A line of work addressed text-to-3D via SDS[48, 34], studied the compositionality of scenes via SDS[33, 13, 17, 43], and more recently via Gaussian Splatting [60]. These methods can create realistic 3D assets from text via a small number of images; however, they are limited in terms of number of objects within a scene while capturing their relationships. Comparably, a set of works used diffusion models to learn 3D shapes explicitly [12, 35]. Finally, diffusion models have recently been shown to be useful for graph generative processes [55, 54, 24]. EchoScene draws from the recent advances in diffusion models, where both shape and layout generation are through graph diffusion processes.
@@ 5.1 Scene Fidelity
5.2 Graph Constraints
We optionally manipulate scene graphs in the latent by either adding a node and relevant edges to the graph or changing relations between two nodes, as illustrated in Fig. 2. As shown in the top and middle row of Tab. 2, for both node addition and relation change, EchoLayout and EchoScene perform better on most metrics. 3D-SLN, CommonScenes, and Graph-to-3D, for instance, loose the constraints smaller/bigger than and close by after graph manipulation, while ours largely keeps it. We believe that the improvement is caused by our scene graph diffusion strategy, where a graph-consistent sample is generated within a realistic appearance distribution.
@@ 5.2 Graph Constraints
The decrease in symmertical category compared with CommonScenes is likely caused by the same effect. Inspection of the data reveals that symmertical is an extremely rare label in the supplementary of CommonScenes [58], with an occurrence of only 0.9%. Thus, the learned part of the distribution is underrepresented and less accurate. The residual echo correction diffuses the latent code towards the learned distribution, resulting in a slight variation. In the bottom row, we report the results without manipulation. All baselines show good performance, yet our EchoLayout and EchoScene are still competitive on their own track. Notably, even though InstructScene shows impressive performance on the global generation fidelity shown in Tab. 1, it struggles to maintain the multiple graph constraints simultaneously, as we described in Sec. 1, while our information echo scheme can sufficiently aggregate relations and condition each process to achieve better results. Distribution-dependant effects from data equally manifest in the results where we can see some drop for symmertical again. Interestingly, the effect is less pronounced for a lower parametric final task such as layout generation. This can be explained by the fact that the variability of bounding boxes is smaller compared to shape variations.
@@ 5.3 Ablation study
We ablate three parts and report the performance in Tab. 4. Firstly, we ablate $\pi(t)$ to check the influence on the model without explicit time information. We observe a marginal decrease, indicating even without $\pi(t)$, the model still learns temporal information through every denoising step. Secondly, we exclude the shape echoes to degrade the shape branch. We clearly see a drop in the fidelity evaluation, where ‘mSG’ focusing on the layouts is not influenced. Thirdly, we substitute cross-attention for concatenation to check if the condition way influences the information echo scheme. We also observe a drop in all metrics.
@@ 5.4 Application
The limitation of EchoScene is that it can generate semantic yet only textureless scenes, which prohibits it from downstream tasks requiring photorealistic textures. Yet, the limitation can be alleviated by directly equipping EchoScene with a subsequent off-the-shelf texture generator. Here, we demonstrate that textured scenes can be generated via SceneTex [9] in Fig. 5.
@@ 6 Conclusion
Acknowledgements
We are grateful to Google University Relationship GCP Credit Program for supporting this work by providing computational resources.
@@ Appendix 0.E Additional Experimental Details
Graph-to-3D series [15]. This series includes one generative method and three object retrieval methods. First, the full generative version Graph-to-3D, stacking two VAE-based branches for object and layout generation, respectively. Second, Graph-to-Box, the single layout branch focusing on the object retrieval task. Third, Progressive, a modified baseline upon Graph-to-Box, specifically adding objects one by one in an autoregressive manner. Fourth, 3D-SLN [36], sharing the same architecture as Graph-to-Box, but without layout standardization during training. We follow the illustration in the supplementary of CommonScenes.
CommonScenes series [58]. This series includes the fully generative version CommonScenes, and its layout branch for object retrieval, CommonLayout. We follow the illustration in the supplementary of CommonScenes.